OpenStudy (anonymous):

4 over 3x + 2 over 3 = 6

OpenStudy (anteater):

4/3x + 2/3 = 6 ?

OpenStudy (anteater):

So, the problem is 4 over 3x plus two-thirds equals 6? Just want to make sure I am reading it correctly.

OpenStudy (anonymous):

yes that is correct

OpenStudy (anteater):

Ok, thanks! There is more than one way you could start on this. One way might be to combine the fractions on the left side of the equation.

OpenStudy (anteater):

But to do that, they'd have to have the same denominator.

OpenStudy (anteater):

So, if you wanted a common denominator for 4/3x and 2/3, what would your common denominator be?

OpenStudy (anonymous):

Isnt the two 3s the common denominator already?

OpenStudy (anteater):

Close! Both denominators have 3 as a factor. :) What you want is the least common multiple of 3 and 3x.

OpenStudy (anonymous):

That would be 1 correct?

OpenStudy (anteater):

Actually, you want an expression that both 3 and 3x would divide into

OpenStudy (anteater):

which would be 3x.

OpenStudy (anteater):

When you said "3" what you gave was the greatest common factor (the largest thing that would divide into both 3 and 3x). We wanted the smallest expression that both 3 and 3x could divide into, which is 3x. Does that make sense?

OpenStudy (anteater):

If not, I will think of another way to word it. :)

OpenStudy (anonymous):

Yes that does, Thank you!

OpenStudy (anteater):

Ok, good :) So 4/3x already has 3x as a denominator. And we want to change 2/3 into an equivalent fraction that has 3x as a denominator also.

OpenStudy (anteater):

So, what can you multiply 2/3 by so that you have 3x in the denominator?

OpenStudy (anteater):

Instead of just 3.

OpenStudy (anonymous):

Im sorry, i have no idea

OpenStudy (anteater):

That's ok. This is what I was getting at: You want both fractions to have the same denominator, so that you can add them together. In order to do that, you have to find a common denominator, which is the least common multiple of the denominators of the fractions you have. In this case, the denominators were 3 and 3x, and so the least common multiple - an expression that both 3 and 3x are factors of - is 3x. Since the 4/3x already has 3x as a denominator, we leave it alone. To change the 2/3 to an equivalent fraction that has 3x as a denominator, we can multiply 2/3 by x/x.

OpenStudy (anteater):

The reason we can multiply by x/x is because x/x = 1.

OpenStudy (anteater):

And when you multiply a number by 1, what do you get?

OpenStudy (anteater):

You get the same number you started out with, right? When you multiply 2/3 by 1, you still have 2/3.

OpenStudy (anteater):

So, since x/x = 1, when you multiply 2/3 by x/x, you get a fraction that still is the same amount as 2/3, but it just looks different: 2x/3x

OpenStudy (anteater):

2x/3x is equivalent to 2/3, but it has the denominator we want. Does that make sense?

OpenStudy (anonymous):

Yes it does!

OpenStudy (anteater):

Ok, good! :D Sorry I am so wordy! ;)

OpenStudy (anteater):

So now we have 4/3x + 2x/3x = 6

OpenStudy (anteater):

And since the fractions have the same denominator, we can combine the sum into one fraction: (4 + 2x)/3x = 6

OpenStudy (anteater):

Do you have an idea of what you would do next?

OpenStudy (anonymous):

Not really, no.

OpenStudy (anteater):